# HERGC: Heterogeneous Experts Representation and Generative Completion for Multimodal Knowledge Graphs

### v1

@@ 3.4 Generative LLM predictor
The Generative LLM Predictor (GLP) aims to predict the correct entity from a set of candidates given a query triple. The query is typically formulated as a question derived from the query entity and relation. To simplify the design, we use an instruction-based prompt that directly asks the LLM to complete an incomplete triple by selecting an appropriate entity. Additionally, we inject the learned fused multimodal embeddings into the LLM to provide auxiliary information for prediction.
Prompt Template. Taking the tail prediction scenario as an example, we first use HERR to retrieve the ranking of all candidates based on the query $(h_{q},r_{q},?)$, ensuring that the resulting triples do not already exist in the MMKG. We then select the top-$k$ candidates, denoted as $C=[e_{1},e_{2},\ldots,e_{k}]$. A natural language question $Q$ is generated based on the query relation $r_{q}$ and entity $h_{q}$. Finally, we construct a prompt $P$ by combining the instruction $I$, the question $Q$, the candidate list $C$, and the fused embeddings $V$ of $h_{q}$ and each $e\in C$:
$P=[I,Q,C,V].$ (11)
Appendix A.2 presents the detailed prompt template.
LoRA Fine-tuning. We fine-tune open-source LLMs with Low-Rank Adaptation (LoRA), using a small number of query–answer pairs. This lightweight adaptation teaches the model to follow our completion instruction while largely relying on its pretrained knowledge. The injected multimodal embeddings also help guide the LLM toward the correct choice, enabling accurate prediction.
@@ 4.1 Experiment Setup
HERGC 38.89 33.12 41.14 47.86 39.82 36.73 41.42 44.84 40.76 33.09 45.07 54.88
@@
Baseline Methods. For MMKGC, we consider both the classic method based on unimodal design and the advanced method based on multimodal design. (1) For unimodal methods, we mainly consider several classic knowledge graph embedding methods: TransE (Bordes et al., 2013), RotatE (Yang et al., 2014), DisMult Sun et al. (2019), GC-OTE (Tang et al., 2019) and TuckER (Balažević et al., 2019). The baseline comparisons in this paper are based on the reported performance values of these methods (2) For the multimodal methods, we selected a series of state-of-the-art multimodal KGE or KGC models: IKRL (Xie et al., 2016), TBKGC (Mousselly-Sergieh et al., 2018), TransAE (Wang et al., 2019b), MMKRL (Lu et al., 2022), RSME (Wang et al., 2021), OTKGE (Cao et al., 2022), IMF (Li et al., 2023), QEB (Lee et al., 2023), VISTA (Lee et al., 2023), MyGO (Zhang et al., 2025a), MoMoK (Zhang et al., 2025b), MCKGC (Gao et al., 2025). The baseline comparisons in this paper are based on the reported performance values of these methods.
@@ 4.2 Main Results
Table 1 presents the main results of our proposed HERGC compared with advacned unimoodal KGC methods and mutilmodal KGC methods. HERGC achieves the best overall performance on MKG-Y, MKG-W, and DB15K across most evaluation metrics. Compared to multimodal baselines, HERGC matches MoMoK for the top MRR on MKG-W and falls marginally behind MMKRL and MCKGC in Hits@10 on MKG-Y, it consistently outperforms all other methods on the remaining metrics across the three datasets. Notably, HERGC boosts Hits@1 by an absolute $3.42\%$, $3.94\%$, and $3.67\%$ on MKG-Y, MKG-W, and DB15K, respectively, over the strongest baseline in each case.
When compared with unimodal KGC models, HERGC demonstrates substantial gains across all metrics. Specifically, relative to TuckER, HERGC improves MRR by $27.04\%$, $7.48\%$, and $20.38\%$ on MKG-Y, MKG-W, and DB15K, respectively, highlighting the effectiveness of incorporating multimodal information for knowledge graph completion.
@@ 4.3 Ablation Studies
To verify the rationality of the HERGC design, we conduct an ablation study consisting of two parts: (1) ablation of modality-specific inputs to assess the contribution of each modality and the model’s ability to leverage multimodal information, and (2) ablation of key components within HERGC, including the design of each part of the retriever and the LLM predictor. The results of the ablation experiment on three datasets are shown in Table 2.
w/o MKG-W MKG-Y DB15K
@@
HERGC (raw) 38.89 33.12 39.82 36.73 40.76 33.09
@@
For For modality ablation, we individually remove the textual, visual, and structural information. In all cases, performance declines, indicating that each modality contributes meaningfully to the model’s predictions and that HERGC effectively integrates multimodal information.
For component ablation, we examine the impact of removing complex PHM experts, the GMU fusion module, and relation-awareness in the retriever, as well as embedding injection in the LLM predictor. Removing any of these components results in performance degradation, highlighting their importance. Notably, omitting the embedding injection also leads to a performance drop, indicating that incorporating exogenous fused multimodal embeddings enriched with graph context indeed enhances the LLM’s reasoning capability.
@@ 4.4 Representation Visualization
We use t-SNE to visualize the entity representations learned by the HERR on DB15K, providing an intuitive view to directly assess its effectiveness. Specifically, we select entities from the following types: "Writer", "Singer", "Flim", "Company", "City", "Language" and "College". We compared the fused multimodal embeddings against individual modality embeddings, each projected from high-dimensional space to two dimensions. The visualization result is shown in Figure 3.
@@
The fused embeddings form almost perfectly separated clusters for each entity type, with clear inter-type boundaries and uniform intra-type distributions. By contrast, image-only embeddings exhibit highly entangled regions; structure-only embeddings fail to distinguish the “Language” cluster and yield a diffuse “Writer” grouping; and text-only embeddings conflate “Writer” and “Singer” entities—likely due to their lexical similarity (e.g., names). These observations confirm that HERR effectively integrates multimodal signals to learn high-quality entity representations.
@@ 4.5 LLM Predictor Exploration
We further investigate the LLM predictor by examining two factors: (1) the effect of varying the candidate set size $k$ on model performance and inference time, and (2) the impact of using LLMs with different parameter sizes.
@@
We also assess the effect of LLM size on the predictor. Table  3 reports the performance and time cost of HERGC using LLMs of different scales. From the results. Although the 3 B model reduces inference time by roughly 30% compared to the 8 B variant, it suffers a modest decline in accuracy, indicating that the more knowledge and better reasoning ability of the larger LLM is indeed helpful for MMKGC prediction. Additional experiment results with large visual models (LVM) and large multi-model models (LMM) are show in Appendix A.5.
Dataset MRR ($\Delta$) Hits@1 ($\Delta$) Hits@3 ($\Delta$) Hits@10 ($\Delta$) Time ($\Delta\%$)
MKG-W 37.02 (-1.87) 31.41 (-1.71) 39.58 (-1.56) 46.71 (-1.15) 3796 (-32.8)
MKG-Y 38.72 (-1.10) 35.43 (-1.30) 40.44 (-0.98) 44.16 (-0.68) 2448 (-32.5)
DB15K 39.61 (-1.15) 31.18 (-1.91) 44.47 (-0.60) 54.34 (-0.54) 5509 (-29.6)
@@ 5 Conclusion
In this paper, we introduce HERGC, a novel generative framework for multimodal knowledge graph completion that adopts a generative paradigm to addresses the limitations of prior MMKGC approaches. HERGC employs a Heterogeneous Experts Representation Retriever to capture modality-specific signals, fuse them into rich entity embeddings, and retrieve a compact candidate set, thereby integrating multimodal context and narrowing the LLM’s search space. It then uses a flexible Generative LLM Predictor to select the correct entity from these candidates, leveraging the LLM’s inherent knowledge and reasoning capabilities to compensate for previous methods’ shortcomings. Extensive experiments on three standard MMKGC benchmarks demonstrate that HERGC achieves state-of-the-art performance while remaining effective and robust. By fully exploiting each modality’s information, the retriever learns high-quality fused multimodal embeddings, and the LLM predictor readily benefits from advances in LLM technology, making HERGC both powerful and easy to implement. Our HERGC bridges the gap between generative paradigm and MMKGC, offering a generalizable framework for future MMKGC research.
6 Limitations
Although HERGC demonstrates strong performance, its effectiveness is inherently bounded by the capabilities of the underlying LLMs. In our additional experiments, the LMM and LVM underperformed, likely due to their potential not being fully explored or insufficient adaptation to the MMKGC task. Moreover, our approach utilizes commonly used encoders, VGG and BERT, to extract coarse-grained multimodal embeddings. While effective, these general-purpose encoders may overlook modality-specific fine-grained features critical for accurate reasoning. Future work could investigate advanced feature extraction techniques tailored to each modality, enabling richer and more discriminative representations. Additionally, scaling HERGC to other MMKG tasks, such as paths reasoning and question answering, and exploring more efficient training paradigms, such as instruction tuning or parameter-efficient adaptation strategies, could further enhance its generalizability and practicality.
@@ A.2 Prompt Template
You are an excellent linguist. The task is to predict the head or tail based on the given incomplete triple, and you only need to answer one entity. The answer must be in (’candidate1’, ’candidate2’, ’candidate3’, ’candidate4’, ’candidate5’, ’candidate6’, ’candidate7’, ’candidate8’, ’candidate9’, ’candidate10’, ’candidate11’, ’candidate12’, ’candidate13’, ’candidate14’, ’candidate15’, ’candidate16’, ’candidate17’, ’candidate18’, ’candidate19’, ’candidate20’).
@@ A.3 Model Training
We train the retriever HERR using the training and test sets following the standard dataset splits of MKG-W, MKG-Y, and DB15K. For training the GLP, we fine-tune the LoRA module with a small number of samples. Specifically, we employ a consistent prompt template to transform the sample triples into query–candidates formats for training. Notably, since the retriever is trained on the training set, the correct entity often receives a high score and is consistently ranked first. To prevent the LLM from overfitting to this shortcut—i.e., learning the retriever’s ranking pattern rather than making predictions based on textual content—we follow previous work (Wei et al., 2023; Liu et al., 2024) and use the validation set to construct the fine-tuning data for the LLM. Concretely, for MKG-W and MKG-Y, we split the original validation set into a training/validation split for LLM fine-tuning at a 9:1 ratio. For DB15K, we randomly sample 5,000 triples from its original validation set and similarly divide them into training and validation subsets using a 9:1 ratio. The test sets remain identical to the original benchmarks, and we perform both head and tail entity prediction for each test triple, in line with standard KGC evaluation protocols.
@@ A.4 Evaluation Metrics
where $N$ is the total number of test queries. MRR ranges from 00 to $1$; higher values indicate better overall ranking quality.
@@
[…]
@@
A.5 Evaluating LVM and LMM as Predictors
To further investigate the GLP component, we experimented with replacing the LLM in GLP with a large vision model (LVM) or a large multimodal model (LMM), enabling the predictor to directly incorporate the image of the query entity in addition to textual input. The results, presented in Table 6, show that this modification did not lead to the expected improvements. Specifically, substituting the LLM with an LVM resulted in a marked reduction in overall performance, whereas replacing it with an LMM offered no substantial benefit, yielding only marginal gains in MRR ($+1.1\%$) and Hits@10 ($+0.8\%$).

### v2

@@ 4.3 Generative LLM predictor
The Generative LLM Predictor (GLP) aims to predict the correct entity from a set of candidates given an incomplete query triple. Each query is reformulated as a natural language question derived from the query entity and relation. We adopt an instruction prompt that directly asks the LLM to complete an incomplete triple by choosing the most suitable entity.
Prompt Template. Taking the tail prediction scenario as an example, we first use HERR to retrieve the ranking of all candidates based on the query $(h_{q},r_{q},?)$, ensuring that the resulting triples do not already exist in the MMKG. We then select the top-$k$ candidates, denoted as $C=[e_{1},e_{2},\ldots,e_{k}]$. A natural language question $Q$ is generated based on the query relation $r_{q}$ and entity $h_{q}$. Finally, we construct a prompt $P$ by combining the instruction $I$, the question $Q$, the candidate list $C$, and the entity descriptions $D$ (including text descriptions for $h_{q}$ and each $e\in C$, and the image for $h_{q}$ when using multimodal LLMs such as LLaMA-3-Vision):
$P=[I,Q,C,D].$ (11)
Using this prompt, closed-source LLMs can perform prediction via in-context learning without additional training.
LoRA Fine-tuning. For open-source LLM, We perform fine-tuning with Low-Rank Adaptation (LoRA) on a small number of query–answer pairs. In this setting, we inject the fused embedding into the LLM via prompt using an adapter layer. Thus, the prompt template becomes:
$P=[I,Q,C,D,E],$ (12)
where $E$ denotes the fused embeddings of $h_{q}$ and each $e\in C$. This lightweight adaptation enables the model to follow our completion instruction while largely relying on its pretrained knowledge. The injected multimodal features provide additional grounding signals, guiding the model toward more accurate predictions.
@@ 5.1 Experiment Setup
HERGCRetriever-only 36.22 30.56 38.32 46.81 38.42 35.11 40.16 44.29 38.76 30.67 42.71 54.20
HERGCGPT-4 38.28 32.03 41.80 47.82 39.23 35.69 42.22 45.09 39.70 31.22 45.09 55.38
HERGCLLaMA-3 39.12 33.65 41.67 48.12 39.82 36.73 41.42 44.84 40.95 33.47 45.66 55.12
HERGCLLaMA-3-Vision 38.76 33.01 41.43 48.54 40.26 36.31 40.91 45.22 40.28 32.30 45.20 54.67
@@
Baseline Methods. For MMKG, we consider both the classic method based on unimodal design and the advanced method based on multimodal design. (1) For unimodal methods, we mainly consider several classic knowledge graph embedding methods: TransE (Bordes et al., 2013), RotatE (Sun et al., 2019), DisMult Yang et al. (2015), ComplEx (Trouillon et al., 2016) and TuckER (Balažević et al., 2019). The baseline comparisons in this paper are based on the reported performance values of these methods (2) For the multimodal methods, we selected a series of powerful multimodal KGE or KGC models: IKRL (Xie et al., 2017), TBKGC (Mousselly-Sergieh et al., 2018), TransAE (Wang et al., 2019b), MMKRL (Lu et al., 2022), RSME (Wang et al., 2021), OTKGE (Cao et al., 2022), IMF (Li et al., 2023), QEB (Lee et al., 2023), VISTA (Lee et al., 2023), MyGO (Zhang et al., 2025a), MoMoK (Zhang et al., 2025b), MCKGC (Gao et al., 2025). The baseline comparisons in this paper are based on the reported performance values of these methods.
@@ 5.2 Main Results
Table 1 presents the main results of our proposed HERGC compared with advanced unimodal and multimodal KGC methods. HERGC consistently achieves the best overall performance on three datasets across most evaluation metrics, demonstrating the effectiveness of its design in leveraging multimodal information and the reasoning capabilities of LLMs. Notably, HERGC improves Hits@1 on MKG-Y, MKG-W, and DB15K by $7.44\%$, $3.94\%$, and $3.37\%$, respectively, over the strongest baseline on each dataset.
We also assess the impact of different LLM predictors within GLP. LLaMA-3, after fused embeddings injection and lightweight LoRA tuning, yields consistently strong results, whereas LLaMA-3-Vision offers only marginal gains, likely because the images do not directly carry the discriminant information of the current relations. Furthermore, comparisons with GPT-4 show that with fused embeddings and lightweight fine-tuning, the open-source model can outperform the powerful closed-source model, indicating that structural reasoning ability can be enhanced through multimodal integration.
@@ 5.3 Ablation Studies
To verify the rationality of the HERGC design, we conduct an ablation study consisting of three parts: (1) ablation of modality-specific inputs to assess the contribution of each modality and the model’s ability to leverage multimodal information; (2) ablation of key components within HERGC, including the design of each part of the retriever and the LLM predictor; and (3) replacement of the default Tucker with alternative score functions (TransE, RotatE, and ComplEx). The results on three datasets are shown in Table 2.
Setting MKG-W MKG-Y DB15K
@@
Score Functions (w/)
TransE 33.27 26.32 34.78 28.80 28.58 20.81
RotatE 34.43 27.12 35.24 31.12 28.12 19.43
ComplEx 27.52 20.03 31.68 25.78 32.26 24.50
HERGCLLaMA-3 39.12 33.65 39.82 36.73 40.95 33.47
@@
For modality ablation, we individually remove the textual, visual, and structural information. In all cases, performance declines, indicating that each modality contributes meaningfully to the model’s predictions and that HERGC effectively integrates multimodal information. For component ablation, we examine the impact of removing complex PHM experts, the RaGMU fusion module, and relation-awareness in the retriever, as well as embedding injection in the LLM predictor. Removing any of these components results in performance degradation, highlighting their importance. Notably, omitting the embedding injection also leads to a performance drop, indicating that incorporating exogenous fused multimodal embeddings enriched with graph context indeed enhances the LLM’s reasoning capability. Furthermore, the comparison of different scoring functions further validates the effectiveness of using TuckER.
@@ 5.4 Representation Visualization
We use t-SNE to visualize the entity representations learned by the HERR on DB15K and compared them against individual modality embeddings, providing an intuitive view to directly assess its effectiveness. We select entities from the following types: "Writer", "Singer", "Flim", "Company", "City", "Language" and "College". As shown in Figure 3, the fused embeddings form almost perfectly separated clusters for each entity type, with clear inter-type boundaries and uniform intra-type distributions. By contrast, image-only embeddings exhibit highly entangled regions; structure-only embeddings fail to distinguish the “Language” cluster and yield a diffuse “Writer” grouping; and text-only embeddings conflate “Writer” and “Singer” entities—likely due to their lexical similarity (e.g., names). These observations confirm that HERR effectively integrates multimodal signals to learn high-quality entity representations.
@@ 5.5 LLM Predictor Exploration
We further investigate the GLP when using open-source LLM by examining two factors: (1) the effect of varying the candidate set size $k$ on model performance and fine-tuning time, and (2) the impact of using LLMs with different parameter sizes.
Table  3 reports the performance and time cost of HERGC using LLMs of different scales. From the results. Although the 3 B model reduces inference time by roughly 30% compared to the 8 B variant, it suffers a modest decline in accuracy, indicating that the more knowledge and better reasoning ability of the larger LLM is indeed helpful for MMKGC prediction.
@@
Dataset MRR ($\Delta$) Hits@1 ($\Delta$) Time ($\Delta\%$)
MKG-W 37.02 (-2.10) 31.41 (-2.24) 3796 (-32.8)
MKG-Y 38.72 (-1.10) 35.43 (-1.30) 2448 (-32.5)
DB15K 39.61 (-1.34) 31.18 (-2.29) 5509 (-29.6)
@@ 6 Conclusion
In this paper, we present HERGC, a novel generative framework for MMKGC. HERGC comprises a Heterogeneous Experts Representation Retriever (HERR), which fuses multimodal signals into high-quality entity embeddings and retrieves a compact candidate set, and a Generative LLM Predictor (GLP), which predicts the correct entity from candidates and supports both open- and closed-source LLMs. Extensive experiments on three standard MMKGC benchmarks demonstrate that HERGC achieves state-of-the-art performance and consistent robustness. HERGC bridges the generative paradigm and MMKGC, providing a generalizable solution for future research.
@@ A.2 Prompt Template
You can refer to the entity descriptions: query entity’: [image], query entity’: [description], ’candidate1’: [description], ’candidate2’: [description], ’candidate3’: [description], ’candidate4’: [Placeholder], ’candidate5’: [description], ’candidate6’: [Placeholder], ’candidate7’: [Placeholder], ’candidate8’: [description], ’candidate9’: [description], ’candidate10’: [description], ’candidate11’: [Placeholder], ’candidate12’: [description], ’candidate13’: [description], ’candidate14’: [description], ’candidate15’: [description], ’candidate16’: [description], ’candidate17’: [description], ’candidate18’: [description], ’candidate19’: [description], ’candidate20’: [description].
@@ A.3 Model Training
We train the retriever HERR using the training and test sets following the standard dataset splits of MKG-W, MKG-Y, and DB15K. For training the GLP when using open-source LLMs, we fine-tune the LoRA module with a small number of samples. Specifically, we employ a consistent prompt template to transform the sample triples into query–candidates formats for training. Notably, since the retriever is trained on the training set, the correct entity often receives a high score and is consistently ranked first. To prevent the LLM from overfitting to this shortcut—i.e., learning the retriever’s ranking pattern rather than making predictions based on textual content—we follow previous work (Wei et al., 2023; Liu et al., 2024) and use the validation set to construct the fine-tuning data for the LLM. Concretely, for MKG-W and MKG-Y, we split the original validation set into a training/validation split for LLM fine-tuning at a 9:1 ratio. For DB15K, we randomly sample 5,000 triples from its original validation set and similarly divide them into training and validation subsets using a 9:1 ratio. The test sets remain identical to the original benchmarks, and we perform both head and tail entity prediction for each test triple, in line with standard KGC evaluation protocols.
For computational efficiency, the addition of fine-tuning LLM does not introduce significant overhead. As shown in Table 6, LLM fine-tuning and inference account for only 8.05%, 7.15%, and 8.42% of the total training time on MKG-W, MKG-Y, and DB15K, respectively. The total training time remains reasonable for a multimodal KGC task of this scale.
Dataset HERR GLP Total
MKG-W 17 h 53 min 1 h 34 min 19 h 27 min
MKG-Y 13 h 51 min 1 h 04 min 14 h 55 min
DB15K 23 h 33 min 2 h 10 min 25 h 43 min
Table 6: Training time breakdown of the HERR and GLP when using LlaMA-3.
@@ A.4 Evaluation Metrics
where $N$ is the total number of test queries. MRR ranges from 0 to $1$; higher values indicate better overall ranking quality.
@@
$\text{Hits@k}=\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}\!\bigl{[}\operatorname{rank}_{i}\leq k\bigr{]},$
@@
A.5 Exploring LVM and LMM as Predictors
To further investigate the GLP component, we experimented with replacing the LLM in GLP with a large vision model (LVM) or a large multimodal model (LMM), enabling the predictor to directly incorporate the image of the query entity in addition to textual input. The results, presented in Table 7, show that this modification did not lead to the expected improvements. Specifically, substituting the LLM with an LVM resulted in a marked reduction in overall performance, whereas replacing it with an LMM offered no substantial benefit, yielding only marginal gains in MRR ($1.1\%$) and Hits@10 ($0.8\%$).